OpenStudy (anonymous):

What is the x-intercept of the graph: 2x + 4y = -10

OpenStudy (anonymous):

when you're looking for an x-intercept, your y is going to equal zero

OpenStudy (anonymous):

The x-intercept happens when y=0. Plug 0 in for y to get 2x+4(0)=-10 and solve for x. That will be your x-intercept.

OpenStudy (anonymous):



OpenStudy (anonymous):

soooo just set y=0 in your equation and solve that shiyat

OpenStudy (zzr0ck3r):

2x+4y=-10 make y = 0 2x + 0 = -10 what is x?